find the values of x and y for which 2y+5= x-1 and y-2=3x is true
Answer by mathbird(71)

2y+5= x-1---------->eq 1
y-2=3x------------->eq 2
from eq 2:y=3x+2
put in eq 1:2(3x+2)=x-1
6x+4=x-1
6x-x=-1-4
5x=-5
x=-1
now put the value in any equation 1 or 2
right now putting x value in eq2
y-2=3x
y=2+3(-1)
y=2-3
y=-1